I need the answers ASAP
Apps Bookmarks ok Psycho-pass oog Tv Goals I Googl
Answer #1

1)

According to Gauss law the electric flux is given by

phi = Qin / epsilon not

phi = 5 * 10^-6 / 8.854 * 10^-12

for each face

phi = 1/6 * 5 * 10^-6 / (8.854 * 10^-12)

phi = 9.4 * 10^4 N-m^2/C


3)

a)

E = sigma / 2*epsilonnot

9.5 * 10^4 = sigma / (2*8.854*10^-12)

sigma = 1.68 *10^-6 C

upper face sigma = 1.68 *10^-6 C


lower face sigma = 1.68 *10^-6 C


b)

charge Q = sigma * A

Q = 1.68 *10^-6 * 0.31^2

Q = 0.16 * 10^-6 C

upper face Q = + 0.16 * 10^-6 C

lower face Q = - 0.16 * 10^-6 C

4)

a) at 5 cm E = 0 ( because no charge is present in side

b)

E = k * Q / r^2

E = 9 * 10^9 * 57 * 10^-6 / 0.44^2

E = 2649793.3 N/c